Determine the maximum error in quadratic interpolation at equispaced points ?
1
Expert's answer
2020-11-03T16:49:13-0500

Let assume P1(y) is the linear polynomial interpolating f (y) at y0 and y1.

If assumed that f (y) is twice continuously differentiable on an interval [m,n] which contains the points y0 <y1.

Then for m ≤ y ≤ n,

f (y)−P1(y) = (y −y0)((y −y1)/2) f0(cy) for some cy between the maximum and minimum of y0, y1, and y.

P1(x) is usually used as an approximation of f (y) for y ∈ [y0,y1].

Then for an error bound, |f (y)−P1(y)|≤ (y −y0)(y1 −y).

Easily, with h =y1 −y0,

max x0≤x≤x1 (y −y0)(y1 −y) = h2/4.

Therefore, |f (y)−P1(y)|≤ h2, y ∈ [y0,y1].